OpenStudy (anonymous):

There are (3^2)^4 ⋅ 3^0 bacteria in a Petri dish. What is the total number of bacteria in the dish?

OpenStudy (anonymous):

@iGreen

OpenStudy (igreen):

Simplify the exponent inside the parenthesis. \(\sf 3^2=~?\)

OpenStudy (anonymous):

9

OpenStudy (igreen):

Correct, now simplify: \(\sf (9)^4\)

OpenStudy (anonymous):

6561?

OpenStudy (igreen):

Yep! And what's \(\sf 3^0\)?

OpenStudy (anonymous):

1

OpenStudy (igreen):

Correct, that leaves us with: \(\sf 6561\times 1\)

OpenStudy (anonymous):

6561

OpenStudy (phi):

they may want the answer as \[ 3^8\] you can find this answer by knowing something to the 4th power means multiply by itself four times, so \[ \left(3^2\right)^4= 3^2\cdot 3^2 \cdot 3^2\cdot 3^2\] you should know \( 3^2= 3 \cdot 3\) so you can also write this as \[ 3^2\cdot 3^2 \cdot 3^2\cdot 3^2= 3 \cdot 3 \cdot 3 \cdot 3 \cdot 3 \cdot 3 \cdot 3 \cdot 3 \] use the "short way" to show you have 3 multiplied by itself 8 times

OpenStudy (phi):

in other words \[ 3 \cdot 3 \cdot 3 \cdot 3 \cdot 3 \cdot 3 \cdot 3 \cdot 3=3^8\]

OpenStudy (anonymous):

okay thank you that is what my question asked for

OpenStudy (igreen):

Yes, @phi is just explaining another way to do it. \(\sf 3^8\) also gives us the answer of 6561

OpenStudy (igreen):

Nice work

OpenStudy (anonymous):

ok thanks